OpenStudy (anonymous):

prove: cosx/1-sinx=tanx +1/cosx

OpenStudy (callisto):

\[LS=\frac{cosx}{1-sinx}\]\[= \frac{cosx(1+sinx)}{(1-sinx)(1+sinx)}\]\[ = \frac{cosx+sinxcosx}{1-sin^2x} \]\[= \frac{cosx+sinxcosx}{cos^2x}\]\[= \frac{cosx}{cos^2x}+\frac{sinxcosx}{cos^2x}\]Can you continue?

OpenStudy (anonymous):

I cant

OpenStudy (callisto):

Do you understand the previous workings?

OpenStudy (anonymous):

thank you callisto I got it thanks

OpenStudy (callisto):

Welcome!~